A jar contains eight marbles. One of them is blue. If five marbles are independently drawn from the jar, what is closest to the probability that the blue marble is drawn twice?
Answer #1

Here we have eight marbles out of which one is blue.

If marbles are drawn without replacement:

Since there are only one blue marble so it is not possible to get blue marble twice therefore

P(getting blue marble twice) = 0

----------------------------------------------------------

If marbles are drawn with replacement:

Then probability of getting blue marble remain same each time and it is equal to p = 1/ 8 = 0.125

Let X is a random variable shows the number of blue marbles out of 5. Here X has binomial distribution with following parameters n=5 and p=0.125.

The probability that the blue marble is drawn twice is

P(X=2)=\binom{5}{2}(0.125)^{2}(1-0.125)^{5-2}=0.1047

-----------------------------